OpenStudy (anonymous):

what is 15% off of $899

jimthompson5910 (jim_thompson5910):

15% of 899 = (15/100) * 899 15% of 899 = (0.15) * 899 15% of 899 = ???

OpenStudy (anonymous):

.15 x 899 then subtract the number you get off 899

jimthompson5910 (jim_thompson5910):

Once you found that result, subtract it from 899